# Quasi-amorphous crystal

### v1

@@ Abstract
The value of the power-law exponent $\epsilon=1.0$ is a signature of an archetypically ’wild’ crystal plasticity in the sense of Weiss et al. (2015). The same value was obtained in discrete dislocation dynamics studies of disorder-free 2D samples containing a fixed number of preexisting dislocations Ispánovity et al. (2014); Salmenjoki et al. (2021) and was also recorded in numerical experiments on pristine crystals using a scalar version of the MTM Zhang et al. (2020). The exponent $\epsilon=1.0$ has been previously interpreted in crystal plasticity framework as representing either dislocation jamming or self-induced glassiness Ovaska et al. (2015); Lehtinen et al. (2016); Zhang et al. (2016).
@@
It is rather striking that exactly the same value of the exponent has been also recorded in the studies of plastic flows in structural glasses Tyukodi et al. (2019); Ferrero and Jagla (2019); Shang et al. (2020). A theoretical understanding of the relation between the emergence of the power law exponent $\epsilon=1.0$ and the structure of the underlying energy landscapes has been developed in the theory of spin glasses. Thus, it was shown that the hierarchical (ultrametric) organization of energy wells in the phase space results in an intermittent, scale-free response to quasistatic deformation  Franz and Spigler (2017); Müller and Wyart (2015); Berthier et al. (2019); Dennis and Corwin (2020). Moreover, it was rigorously proved that in the mean-field limit the associated distribution of energy avalanches must be of a power law with exponent $\epsilon=1.0$. In the spin glass context it has been also understood that the reason behind this particular value of the exponent is the marginal stability of the system  Franz and Spigler (2017); Pázmándi et al. (1999); Doussal et al. (2012).
@@
Since glasses can be viewed as liquids at a quenching threshold, their marginal stability is similar to marginality exhibited by granular matter at a jamming point. If such systems are mechanically driven, the proximity to unstable modes leads to the mixing of statics (stability) and dynamics (instability) making the mechanical response inherently intermittent  Lamp et al. (2022); Rossi and Tarjus (2022). Similarly, since our ’quasi-amorphous crystal’ emerged from arrested dynamics, one can argue that the associated self-generated disorder brought the system from a solid to an effectively glassy state. Furthermore, our results suggest that the implied marginality is not affected by the quasi-brittle yield, which apparently does not compromise the global organization of the energy landscape while, of course, affecting the reachability of its different subdomains Regev and Lookman (2017a, b); Regev et al. (2021).
@@
Note next that the pre-yield value of the cutoff parameter $\lambda=0.0230$ is drastically different from its post-yield value $\lambda=0.0076$. The approximately three-fold decrease of the value of $\lambda$ indicates that large energy drops become substantially more probable in the post-yield regime  Lin et al. (2014); Ruscher and Rottler (2021). The emergent collective behavior is reflected in a much broader extent of avalanche activity which reaches the size of the system suggesting the divergence of the characteristic length $l\sim 1/\lambda$. The physical nature of the implied criticality is presently actively debated Grinstein (1995); Zippelius et al. (2023); Charbonneau et al. (2023); Oyama et al. (2023); Xing et al. (2024); Zhang et al. (2020).
@@
Finally we mention that the observed similarity between the structure of intermittent plastic fluctuations in amorphous and crystalline solids is supported by the fact that elasto-plastic models, typically used to model amorphous plasticity Nicolas et al. (2018); Rossi et al. (2022); Ferrero and Jagla (2019); Fernández Castellanos et al. (2021), are very similar to the MTM which operates within basically the same finite element setting. The difference is that the phenomenological yield thresholds of elasto-plastic models are replaced in the MTM by elastic instabilities originating from the non-convexity of the energy landscape. Accordingly, the fixed linear elastic propagators are replaced by the solution ’on the fly’ of the corresponding nonlinear elasticity problems. It is important that the crucial nonlinearity in the MTM is of universal geometrical nature as it originates from the exact description of finite elastic deformations.
@@
To conclude, we showed that mechanically driven perfect crystals can exhibit quasi-brittle plastic yielding which is remarkably similar to the behavior of well-annealed glassy materials. The implied parallel features of the mechanical response emerge after pristine crystals acquire self-generated disorder resulting from massive dislocation nucleation during the breakdown of an affine elastic configuration. The ensuing quasi-amorphous crystals exhibit pre- and post-yield avalanches with power law statistics whose matching exponents are indicative of the marginality of the associated mechanical system. Adaptation of the same model to realistic 3D crystals will allow one to distinguish between edge and screw dislocations opening the way to the adequate account of such physically important effects as dislocation climb, cross slip and forest hardening.
@@
Acknowledgments. The authors thank N. Gorbushin, S. Patinet, G. Tejedor for helpful discussions. The work of O.U.S. was supported by the grants ANR–21-CE08-MESOCRYSP, ANR-20-CE91-0010, ANR-24-CE91-0002. L. T. was supported by the grants ANR–21-CE08-MESOCRYSP and ERC-H2020-MSCA-RISE-2020-101008140.

### v2

@@ Abstract
The value of the power-law exponent $\epsilon\approx 1.0$ is a signature of an archetypically ’wild’ crystal plasticity in the sense of Weiss et al. (2015). The same value was obtained in discrete dislocation dynamics studies of disorder-free 2D samples containing a fixed number of preexisting dislocations Ispánovity et al. (2014); Salmenjoki et al. (2021) and was also recorded in numerical experiments on pristine crystals using a scalar version of the MTM Zhang et al. (2020). The exponent $\epsilon\approx 1.0$ has been previously interpreted in crystal plasticity framework as representing either dislocation jamming or self-induced glassiness Ovaska et al. (2015); Lehtinen et al. (2016); Zhang et al. (2016).
@@
It is rather striking that exactly the same value of the exponent has been also recorded in the studies of plastic flows in structural glasses Tyukodi et al. (2019); Ferrero and Jagla (2019); Shang et al. (2020). A theoretical understanding of the relation between the emergence of the power law exponent $\epsilon\approx 1.0$ and the structure of the underlying energy landscapes has been developed in the theory of spin glasses. Thus, it was shown that the hierarchical (ultrametric) organization of energy wells in the phase space results in an intermittent, scale-free response to quasistatic deformation  Franz and Spigler (2017); Müller and Wyart (2015); Berthier et al. (2019); Dennis and Corwin (2020). Moreover, it was rigorously proved that in the mean-field limit the associated distribution of energy avalanches must be of a power law with exponent $\epsilon\approx 1.0$. In the spin glass context it has been also understood that the reason behind this particular value of the exponent is the marginal stability of the system  Franz and Spigler (2017); Pázmándi et al. (1999); Doussal et al. (2012).
@@
Since glasses can be viewed as liquids at a quenching threshold, their marginal stability is similar to marginality exhibited by granular matter at a jamming point. If such systems are mechanically driven, the proximity to unstable modes leads to the mixing of statics (stability) and dynamics (instability) making the mechanical response inherently intermittent  Lamp et al. (2022); Rossi and Tarjus (2022). Similarly, since our ’quasi-amorphous crystal’ emerged from arrested dynamics, one can argue that the associated self-generated disorder brought the system from a solid to an effectively glassy state. Furthermore, our results suggest that the implied marginality is not affected by the quasi-brittle yield, which apparently does not compromise the global organization of the energy landscape while, of course, affecting the reachability of its different subdomains Regev et al. (2021).
The situation is not exactly the same for the exponent $\tau$ characterizing the power law statistics of the stress drops $\Delta\Sigma$. Note first that in the pre-yield regime, the power-law exponents for stress and energy are rather similar, $\tau\approx\epsilon$, which suggests that the dissipated energies and the stress drops are roughly proportional, $\Delta W\sim\Delta\Sigma$. Such linear scaling implies some kind of caging, which means a highly constrained regime of dislocation motion with individual avalanches arrested by a regular arrays of obstacles. This may be also the reason behind the strong hardening behavior in the pre-yield regime.
Instead, in the post-yield regime the recovered exponents $\tau$ and $\epsilon$ are markedly different. Here individual avalanches are no longer hindered by local structural constraints and one can assume that the stress drop is only constrained by the corresponding strain increment due to the constraints of (nonlinear) elasticity. Then, if we assume that $\Delta W\sim\Delta\Sigma^{\gamma}$ we can expect that $\epsilon-1=(\tau-1)/\gamma$. In particular, if elasticity is linear, $\gamma=2$, and the measured value is $\tau\approx 1.2$, we can expect that $\epsilon\approx 1.10$ which is in reasonable agreement with our measured value $\epsilon\approx 1.02$ given that the uncertainty for our relatively small system is of the order $\sim\pm 0.2$, see SM for details.
@@
Note next that the pre-yield values of the cutoff parameters $L^{D_{x}}$ are much smaller than their post-yield values which indicates that larger avalanches become more probable  Lin et al. (2014); Ruscher and Rottler (2021). The emergent collective behavior, reflected in such broader extent of avalanche activity, reaches in post yield regime the size of the system suggesting the divergence of the characteristic length. Further differences between pre-yield and post-yield mechanical responses emerge from the analysis of the values of the exponents $D_{x}$ providing information about the morphology of the plastically deforming regions. Thus, in In the pre-yield regime the energy avalanches are characterized by a low fractal dimension $D_{\epsilon}\approx 0.36$, pointing towards highly localized, spatially scattered, macroscopically isolated rearrangements. Instead, in the post-yield regime we obtain the value $D_{\epsilon}\approx 0.98$ which suggests the emergence of 1D-like system spanning bands. The latter reflect highly cooperative nature of the underlying plastic deformation and the physical nature of the implied criticality is presently actively debated Grinstein (1995); Zippelius et al. (2023); Oyama et al. (2023); Xing et al. (2024); Zhang et al. (2020).
@@
Finally we mention that the observed similarity between the structure of intermittent plastic fluctuations in amorphous and crystalline solids is supported by the fact that elasto-plastic models, typically used to model amorphous plasticity Nicolas et al. (2018); Rossi et al. (2022); Ferrero and Jagla (2019); Fernández Castellanos et al. (2021), are very similar to the MTM which operates within basically the same finite element setting. The difference is that the phenomenological yield thresholds of elasto-plastic models are replaced in the MTM by elastic instabilities originating from the non-convexity of the energy landscape. Accordingly, the fixed linear elastic propagators are replaced by the solution ’on the fly’ of the corresponding nonlinear elasticity problems. Note also that the crucial nonlinearity in the MTM is of universal geometrical nature as it originates from the exact description of finite elastic deformations. However, the observed general agreement in the values of the computed exponents suggests that we are dealing with the same universality class.
@@
To conclude, we showed that mechanically driven perfect crystals can exhibit quasi-brittle plastic yielding which is remarkably similar to the behavior of well-annealed glassy materials. The implied parallel features of the mechanical response emerge after pristine crystals acquire self-generated disorder resulting from massive dislocation nucleation during the breakdown of an affine elastic configuration. The ensuing quasi-amorphous crystals exhibit pre- and post-yield avalanches with power law statistics whose matching exponents are indicative of the marginality of the associated mechanical system. Adaptation of the same model to realistic 3D crystals will allow one to distinguish between edge and screw dislocations opening the way to the adequate account of such physically important effects as dislocation climb, cross slip and forest hardening. To be adequate such an analysis should be, of course, performed with the account of the effects of finite temperature Zhang et al. (2023); Li et al. (2024).
@@
Acknowledgments. The authors thank N. Gorbushin for helpful comments. The work of O.U.S. was supported by the grants ANR–21-CE08-MESOCRYSP, ANR-20-CE91-0010, ANR-24-CE91-0002. L. T. was supported by the grants ANR–21-CE08-MESOCRYSP and ERC-H2020-MSCA-RISE-2020-101008140.